# Enhancing Demand Prediction in Open Systems by Cartogram-aided Deep Learning

### v1

@@ III.2 Prediction performance
Figure 6: The prediction result before and after using the cartogram– for year 2019. (a) Scatter plot of $e^{\rm RMSE}_{i}$ for the rental for all cells, before and after applying the cartogram. The straight line stands for $y=x$. (b) The average errors $e^{{\rm RMSE}}$ and $e^{{\rm MAE}}$ for the rental and return. (c) The average Pearson correlation between empirical $X_{i}(t)$ and prediction $\hat{X}_{i}(t)$ for every cell $i$. The height of the bar corresponds to the average value, and the error bar is a standard deviation.
We evaluate the prediction performance of the ST-CGA model for year 2019 by the following errors; the root-mean-square error $e^{\rm RMSE}_{i}$ and the mean absolute error $e^{\rm MAE}_{i}$ for a cell $i$ computed as
@@
with $T=|\mathbb{T}|$. To confirm the effect of the cartogram in Sec. II.3, we compare the errors in Eq. (III.2) before and after applying the cartogram. As a representative case, we exhibit the scatter plot of $e^{\rm RMSE}_{i}$ for the rental in Fig. 6(a), and one sees entirely the lower errors after applying the cartogram. The average errors in Fig. 6(a) are shown as the leftmost bars in Fig. 6(b), and the averages for the other cases (rental/return and error types) are illustrated as well. Using the cartogram for both rental and return cases results in more enhanced prediction than before applying the cartogram, supported by the about 1.6 times lower average errors: for a year, $e_{{\rm RMSE}}$ ($e_{{\rm MAE}}$) estimates the deviation from the empirical data with about eight (five) bicycles not using the cartogram and with about five (three) bicycles using cartogram. We also measure the temporal Pearson correlation between empirical $X_{i}(t)$ and predicted $\hat{X}_{i}(t)$, and the correlation with the slightly higher average and reduced deviation tells us that our method grasps the temporal trend well. Figures 6(b) and 6(c) allow us to conclude that the use of the cartogram improves the prediction regardless of the type of demand (rental or return).
@@ III.3 Prediction for a new station
We can obtain all $\hat{X}_{i}(t)$’s for all cells due to the cartogram without any exception, which allows us to guess the $\hat{x}_{j}(t)$ for a brand-new station $i$ that was not trained before. Not only it opens up the possibility of the prediction for a brand-new station, but also enables the long-period, yearly prediction. The previous studies Zhang et al. (2017); Kim et al. (2019); Li et al. (2019) exclude the brand-new demand caused by the newly installed stations, which results in the prediction in the limited period (a few weeks or months) to avoid the appearance of the new stations. At this point, we would like to emphasize that we surmount both of the short-period prediction and the exclusion of the new stations, by introducing the cartogram.
The estimation results for a new station, labeled as $j=972$, are shown in Fig. 7, for a week representatively. This station was newly installed and solely located in the cell marked by a red arrow in Fig. 1(b) in 2019, so the cell in 2018 had no station and is shown as being empty in Fig. 1(a). The prediction $\hat{x}_{972}(t)$ does not perfectly match the empirical value $x_{972}(t)$ but their maximal difference is less than one bicycle. Moreover, the quasiperiodic behavior shown in the empirical data is well reproduced in the prediction.
@@ IV Conclusion
We have studied how to enhance the prediction of temporal patterns of rentals and return (demand) for public bicycles, as a representative open system. We have adopted the spatial-temporal convolution graph attention (ST-CGA) network as a prediction model, considering long-range interaction between distant stations. To facilitate new demand in brand-new stations and enhance the prediction performance, we have proposed and introduced the cartogram by iterating the Voronoi tessellation and modified the ST-CGA model by exploiting the batch attention and the update method of the node features in the graph attention. Due to the cartogram, we have achieved not only the original goal (prediction of new demand) but also long-period prediction (Fig. 7), and both have not been achieved in previous studies. The logic behind using a cartogram is to infer the unknown data by the average trend of the adjacent data in a cell, similar to the mean-field approach. In addition, our attempt to introduce batch attention instead of self-attention to the ST-CGA network has demonstrated improved performance. Using the batch attention means that the temporal correlation as well as the spatial correlation also plays a crucial role in accurate prediction.
@@
The simple coarse-graining method shares the same spirit as the one with the cartogram, in a sense to infer an unknown value by the average value of neighboring ones. As a reason why using the cartogram outperforms simple coarse-graining despite the similar sense, we conjecture that our cartogram ensures augmentation of training data points, i.e., the number of cells, with the qualitative level maintained; the adjacent stations having similar demand patterns spread more uniformly, which results in the augmentation of the trainable data points (cells), and the statistics of correlation and dispersion within a cell are maintained (Fig. 4), although the variation of the number of stations can be enough to affect the statistics. Not only our suggested iteration method but also other spreading methods to keep or even strengthen the correlation and dispersion could work on predictability. In addition, While we adopt the demand of public bicycles as a representative of open systems, our framework with such cartogram considered may be applied to other open systems in general, to predict their spatio-temporal patterns.
@@ Appendix A The Self-Attention
For a given temporal resolution $r$, we compose the input data set $\mathbb{X}^{r}(t)$ in Eq. (3) based on the $\tau_{r}$-dimensional vector $\mathbf{X}_{i}^{r}(t;\tau_{r})$ in Eq. (2). In the self- or batch attention, it is a first step to embed the $\tau_{r}$-dimensional vector in a hyperdimension $L$, so we firstly obtain a matrix $\mathbf{X}_{\rm emb}^{r}\in\mathbb{R}^{MN\times B\times L}$ in the embedding space ($\mathbb{R}$ being a set of real number). The embedding dimension $L$ is a hyperparameter, and we heuristically choose $d=32$. The $u$th row of $\mathbf{X}_{\rm emb}^{r}$ contains $d$ features of node $i$ at pivot time $t_{b}$ ($u=i+MNb$) but it is impossible to interpret the physical meaning of every feature in machine learnings.

### v2

@@ III.2 Prediction performance
Figure 6: The prediction result before and after using the cartogram– for the year 2019. (a) Scatter plot of $e^{\rm RMSE}_{i}$ for the rental for all cells, before and after applying the cartogram. The straight line stands for $y=x$. (b) The average errors $e^{{\rm RMSE}}$ and $e^{{\rm MAE}}$ for the rental and return. (c) The average Pearson correlation between empirical $X_{i}(t)$ and prediction $\hat{X}_{i}(t)$ for every cell $i$. The height of the bar corresponds to the average value, and the error bar is a standard deviation.
We evaluate the prediction performance of the ST-CGA model for the year 2019 by the following errors; the root-mean-square error $e^{\rm RMSE}_{i}$ and the mean absolute error $e^{\rm MAE}_{i}$ for a cell $i$ computed as
@@
with $T=|\mathbb{T}|$. To confirm the effect of the cartogram in Sec. II.3, we compare the errors in Eq. (9) before and after applying the cartogram. As a representative case, we exhibit the scatter plot of $e^{\rm RMSE}_{i}$ for the rental in Fig. 6(a), and one sees entirely the lower errors after applying the cartogram. The average errors in Fig. 6(a) are shown as the leftmost bars in Fig. 6(b), and the averages for the other cases (rental/return and error types) are illustrated as well. Using the cartogram for both rental and return cases results in more enhanced prediction than before applying the cartogram, supported by the about 1.6 times lower average errors: for a year, $e_{{\rm RMSE}}$ ($e_{{\rm MAE}}$) estimates the deviation from the empirical data with about eight (five) bicycles not using the cartogram and with about five (three) bicycles using cartogram. The cause for higher performance is expected to be the augmentation of the input data $X_{i}$ under the statistical quality control. We also measure the temporal Pearson correlation between empirical $X_{i}(t)$ and predicted $\hat{X}_{i}(t)$, and the correlation with the slightly higher average and reduced deviation tells us that our method grasps the temporal trend well. Figures 6(b) and 6(c) allow us to conclude that the use of the cartogram improves the prediction regardless of the type of demand (rental or return).
@@ III.3 Prediction for a new station
We can obtain all $\hat{X}_{i}(t)$’s for all cells due to the cartogram without any exception, which allows us to guess the $\hat{x}_{j}(t)$ for a brand-new station $i$ that was not trained before. It opens up the possibility of the prediction for a brand-new station and enables the long-period, yearly prediction. The previous studies Zhang et al. (2017); Kim et al. (2019); Li et al. (2019) exclude the brand-new demand caused by the newly installed stations, which results in the prediction in the limited period (a few weeks or months) to avoid the appearance of the new stations. At this point, we would like to emphasize that we surmount both of the short-period prediction and the exclusion of the new stations, by introducing the cartogram.
The estimation results for a new station, labelled as $j=972$, are shown in Fig. 7, for a week representatively. This station was newly installed and solely located in the cell marked by a red arrow in Fig. 1(b) in 2019, so the cell in 2018 had no station and is shown as being empty in Fig. 1(a). The prediction $\hat{x}_{972}(t)$ does not perfectly match the empirical value $x_{972}(t)$ but their maximal difference is less than one bicycle. Moreover, the quasiperiodic behavior shown in the empirical data is well reproduced in the prediction. The extent of predicting new demand that was not trained can be an indicator to assess the model’s capability, beyond the prediction error Sutton et al. (2020).
@@ IV Conclusion
We have studied how to enhance the prediction of temporal patterns of rentals and return (demand) for public bicycles, as a representative open system. We have adopted the spatial-temporal convolution graph attention (ST-CGA) network as a prediction model, considering long-range interaction between distant stations. To facilitate new demand in brand-new stations and enhance the prediction performance, we have proposed and introduced the cartogram by iterating the Voronoi tessellation and modified the ST-CGA model by exploiting the batch attention and the update method of the node features in the graph attention. Due to the cartogram, we have achieved not only the original goal (prediction of new demand) but also long-period prediction (Fig. 7), and both have not been achieved in previous studies. The logic behind using a cartogram is to infer the unknown data by the average trend of the adjacent data in a cell, similar to the mean-field approach. In addition, our attempt to introduce batch attention instead of self-attention to the ST-CGA network has demonstrated improved performance. Using batch attention means that the temporal correlation as well as the spatial correlation also plays a crucial role in accurate prediction.
The fully connected (weighted) network considered herein may sometimes contain redundant information, which can impede higher accuracy or result in higher computational costs. Extracting sparse network structures requires appropriate filtering methods. Global thresholding is one basic method, but it comes with the challenge of selecting the threshold value, which could be addressed by threshold-free methods in machine learning Ziletti et al. (2018). When constructing sparse networks from the outset, understanding important regional features and direct relationships among regions is necessary, which is often difficult to achieve. In such cases, domain-knowledge-free preprocessing can be applied to construct the initial network Jha et al. (2018). The reliability of such networks can be verified using graph neural networks or node2vec Ribeiro et al. (2023); Lopes et al. (2022).
@@
The simple coarse-graining method shares the same spirit as the one with the cartogram, in the sense of inferring an unknown value by the average value of neighboring ones. As a reason why using the cartogram outperforms simple coarse-graining despite the similar sense, we conjecture that our cartogram ensures augmentation of training data points, i.e., the number of cells, with the qualitative level maintained. As seen in Fig. 4, the input coarse-grained data $X_{i}$ after spreading out the stations significantly increases, i.e., by 100 data points. Furthermore, the adjacent stations not only gradually spread more uniformly but also focus on the spatial nearness of the stations possessing similar demand patterns. The deliberate rearrangement method results in the significant augmentation of the trainable data points (cells) with the statistics of correlation and dispersion within a cell maintained (Fig. 4), although the variation of the number of stations can be enough to affect the statistics. Not only our suggested iteration method but also other spreading methods to keep or even strengthen the correlation and dispersion could work better on predictability. In addition, our framework does not require any system-specific features but spatio-temporal input data. Hence, while we adopt the demand of public bicycles as a representative of open systems, our framework with such cartogram considered may be applied to other open systems in general, such as $e$-scooter, taxi allocation, and the incident cases of disease spreading with migration effect Armbruster and Reinert .
@@ Appendix A The Self-Attention
For a given temporal resolution $r$, we compose the input data set $\mathbb{X}^{r}(t)$ in Eq. (3) based on the $\tau_{r}$-dimensional vector $\mathbf{X}_{i}^{r}(t;\tau_{r})$ in Eq. (2). In the self- or batch attention, it is a first step to embed the $\tau_{r}$-dimensional vector in a hyperdimension $L$, so we firstly obtain a matrix $\mathbf{X}_{\rm emb}^{r}\in\mathbb{R}^{MN\times B\times L}$ in the embedding space ($\mathbb{R}$ being a set of real number). The embedding dimension $L$ is a hyperparameter, and we heuristically choose $L=32$. The $u$th row of $\mathbf{X}_{\rm emb}^{r}$ contains $d$ features of node $i$ at pivot time $t_{b}$ ($u=i+MNb$) but it is impossible to interpret the physical meaning of every feature in machine learnings.